OpenStudy (anonymous):

Which expressions show how 5 • 37 can be rewritten using the distributive property? Choose all answers that are correct. A. 5 • 30 + 7 B. 5 • 30 + 5 • 7 C. 5 • 40 – 5 • 7 D. 5 • 40 – 5 • 3

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@charko

OpenStudy (anonymous):

any idea jhonyy9?

OpenStudy (anonymous):

@jhonyy9

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thx so much i appreciate it!!!! ;D

OpenStudy (anonymous):

But also it has to b emore than one answer ;3

OpenStudy (anonymous):

lulz its ok though

OpenStudy (anonymous):

not excatly i have had qs like this before where there was only 1 answer sometimes they try and trick u by the question wording

OpenStudy (anonymous):

i only have 5 of these questions and this is #2

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

;D

OpenStudy (anonymous):

Simplify: 6y + 5z + 8y – 3z A. 11y + 5z B. 14y + 2z C. 14y + 8z D. 16yz

OpenStudy (anonymous):

any one

OpenStudy (anonymous):

@Joel_the_boss

OpenStudy (anonymous):

@charko

OpenStudy (anonymous):

nvm

OpenStudy (anonymous):

Which are the like terms in 2y3 – 4y2 + y + y3? A. They are all like terms. B. 2y3and y3 C. 4y2and y D. 2y3, –4y2, and y3

OpenStudy (anonymous):

guys

OpenStudy (anonymous):

D?

OpenStudy (anonymous):

2y3and y3

OpenStudy (anonymous):

as long as you help

OpenStudy (anonymous):

thank you so much ;D

OpenStudy (anonymous):

Simplify: 5x3 – 2x2 + x + 4x3 A. 8x3 B. 9x6– 2x2+ x C. 9x3– 2x2+ x D. 5x3– 2x2+ x + 4x3 (the expression cannot be simplified)

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

c

OpenStudy (anonymous):

thx :D

OpenStudy (anonymous):

i made an 80%

OpenStudy (anonymous):

i past

jhonyy9 (jhonyy9):

B is right too because 5*37 = 5*30 +5*7=185=150+35=185